# You Only Sample Once: Taming One-Step Text-to-Image Synthesis by Self-Cooperative Diffusion GANs

### v4

@@ Abstract
We introduce YOSO, a novel generative model designed for rapid, scalable, and high-fidelity one-step image synthesis. YOSO integrates the diffusion process with GANs to achieve the best of two worlds. Specifically, we smooth the distribution by the denoising generator itself, performing self-cooperative learning. We show that our method can serve as a one-step generation model training from scratch with competitive performance. Moreover, we show that our method can be extended to finetune pre-trained text-to-image diffusion for high-quality one-step text-to-image synthesis even with LoRA fine-tuning. In particular, we provide the first diffusion transformer that can generate images in one step trained on 512 resolution, with the capability of adapting to 1024 resolution without extra explicit training.
@@ 1 Introduction
Diffusion models (DMs) [46, 12, 50] have recently emerged as a powerful class of generative models, demonstrating state-of-the-art results in many generative modeling tasks, such as text-to-image [40, 60, 6, 7], text-to-video [4, 13], image editing [9, 5, 32] and controlled generation [64, 33]. However, the generation process of DMs requires iterative denoising, leading to slow generation speed. The sluggish generation time, coupled with the intensive computational requirements of large-scale DMs, constitutes a substantial barrier to their practical application and wider adoption.
@@
Sampling from DMs can be regarded as solving a probability flow ordinary differential equation (PF-ODE) [50]. Some previous works [47, 27, 28, 2] focus on developing advanced ODE-solver, for reducing the sampling steps. However, they still require 20+ steps to achieve high-quality generation. Another line is distilling from pre-trained PF-ODEs [49, 26, 29, 30, 41], aiming to predict multi-step solution of PF-ODE solver by one step. Existing methods [29, 30] can achieve reasonable sample quality with 4+ steps. However, it is still challenging to generate high-quality samples with one step.
@@
In contrast, generative adversarial networks (GANs) [8, 38] are naturally built on one-step generation with fast sampling speed. However, it is hard to extend GANs on large-scale datasets due to the serious training challenges, with worse sample quality compared to DMs [42, 17]. In this work, we present a novel approach to combine diffusion process and GANs. Previous works [54, 59, 60] have developed some variants of diffusion-GAN hybrid models. However, they rely on adding noise to smooth the adversarial divergence, yielding indirect learning in the one-step generation. To directly learn the one-step generation by adversarial divergence and to avoid unstable learning, we propose defining the one-step generation distribution on different levels of corrupted samples by diffusion process, where denotes the distribution of noisy data and denotes the denoised distribution parameterized by denoising generator. Then, we regard the one-step generation based on less corrupted samples as ground truth to perform adversarial divergence. This approach can be regarded as a self-cooperative process, as the generator learns from itself. Such an innovative design enables stable training and effective learning for one-step generation. Hence, we name our model YOSO representing You Only Sample Once. Moreover, we extend our YOSO to one-step text-to-image generation based on pre-trained models by several effective training techniques (i.e., latent perceptual loss and latent discriminator for efficient training along with the latent DMs; the informative prior initialization (IPI), and the quick adaption stage for fixing the noise scheduler).
Thanks to our effective designs, we can efficiently and effectively fine-tune existing pre-trained text-to-image DMs (i.e., stable diffusion [40] and PixArt- [6]) for high-quality one-step generation (See Fig. 1). Moreover, we are the first to support Low Rank Adaptation (LoRA) [16] fine-tuning for one-step generation of text-to-image.
@@
We introduce YOSO, a novel generative model that can generate high-quality images with one-step inference.
We propose several principled and effective training techniques enabling the fine-tuning of pre-trained text-to-image DMs for one-step text-to-image, requiring only ~10 A800 days.
We develop the first one-step diffusion transformer [36] with PixArt-, and the first LoRA enabling one-step generation, both of which will be released soon.
We conduct extensive experiments to demonstrate the effectiveness of the proposed YOSO, including image generation from scratch, text-to-image generation fine-tuning, compatibility with existing image customization, and image controllable modules.
2 Related Works
Recently, several works have been developed for few-step text-to-image generation. LCM [29] adapts consistency distillation [49] for stable diffusion (SD) and enables image generation in 4 steps with acceptable quality. InstaFlow [26] adapts rectified flows [25, 24] for SD, facilitating the generation of text-to-image in a mere single step. Despite this, the fidelity of the generated images is still poor. DMD [62] adapts variation score distillation [52] for SD, which requires training an additional DMs for the generator distribution, akin to the discriminator in GANs. Recently, combining DMs and GANs for one-step text-to-image generation has been explored. UFOGen [60] extends DDGANs [54] and SSIDMs [59] to SD by modifying the computation of reconstruction loss from corrupted samples to clean samples. However, it still performs adversarial matching using corrupted samples. ADD [44] introduces a one-step text-to-image generation based on SD. It follows earlier research [43] by employing a pre-trained image encoder, DINOv2 [34], as the backbone of the discriminator to accelerate the training. However, the discriminator design moves the training from latent space to pixel space, which substantially heightens computational demands. Moreover, they directly perform adversarial matches at clean real data, increasing the challenge of training. This requires the need for a better but expensive discriminator design and expensive R1 regularization to stabilize the training. In contrast, we perform the adversarial training by replacing real data with self-generated data to smooth the ground truth. Moreover, compared with UFOGen and ADD, our approaches can be trained from scratch to perform one-step generation, which is not demonstrated by them. Additionally, we extend our method not only to SD but also to PixArt- [6] which is based on diffusion transformer [36]. This demonstrates the wide application of our proposed YOSO.
@@ 3 Background
Diffusion Models. Diffusion models (DMs) [46, 12] define a forward process that gradually transforms samples from data distribution to Gaussian distribution by adding noise in steps with variance schedule : . The corrupted samples can be directly obtained by , where and . The parameterized reversed diffusion process is defined to gradually denoise: . The model can be trained by minimizing the negative ELBO [12, 20]: where is Gaussian posterior derived in [12]. A key assumption in diffusion is that the denoising step size from to is sufficiently small. This assumption ensures the true follows Gaussian distribution, enabling the effectiveness of modeling with Gaussian distribution.
Diffusion-GAN Hybrids. An issue in DMs is that the true does not follow Gaussian distribution when the denoising step size is not sufficiently small. Therefore, in order to enable large denoising step size, Diffusion GANs [54] propose to minimize the adversarial divergence between model and : , where and is imposed by a GAN generator. The capability of a GAN-based formulation, enables much larger denoising step sizes (i.e., 4 steps).
@@ 4 Method: Self-Cooperative Diffusion GANs
4.1 Revisiting of Diffusion GANs
A key issue in Diffusion-GAN hybrid models is that they match the generator distribution at corrupted data distribution. This formulation only indirectly learns the and , which are the distributions used for one-step generation, making the learning process less effective.
@@ 4.2 Our Design
where and . The optimization objective is constructed by the combination of adversarial divergence and KL divergence. Specifically, the adversarial divergence focuses on matching over the distribution level, ensuring the generation quality, and the KL divergence focuses on matching over point level, ensuring the mode cover.
However, it is hard to directly learn the adversarial divergence over clean data distribution, akin to the challenges with GAN training. To tackle these challenges, previous Diffusion GANs [54, 59] have pivoted towards learning the adversarial divergence over the corrupted data distribution. Unfortunately, as analyzed before, such an approach fails to directly match , curtailing the efficacy of one-step generation. Moreover, it also compels the discriminator to fit different levels of noise, leading to limited capability.
@@
Hence given the generator fixed, if we increase the information in , it is supposed that we can get a better distribution. In other words, it is highly likely that is superior to . Motivated by the cooperative approach [55, 56, 57, 11] which uses an MCMC revised version of model distribution to learn the generator, we suggest to use as ground truth to learn , which constructs the following training objective:
where denotes stop-gradient operator. This training objective can be regarded as a self-cooperative approach, since the ‘revised’ samples from and samples from are generated by the same network . Note that we only replace data distribution as in the adversarial divergence for smoothing the learning objective, as recent work [31] has found that it is beneficial to learn the generator with a mix of real data and revised data.
@@
where is the discriminator network. We find that the self-cooperative approach is connected with Consistency Training [49]. However, Consistency Training considers the as an approximated ODE solution of to perform a point-to-point match. In contrast, our proposed objective matches and in marginal distribution level, which avoids the approximated error of ODE.
To further ensure the mode cover of the proposed model, we can add consistency loss to our objective as regularization, which constructs our final training objective:
@@ 5.1 Using pre-trained models for Training
Latent Perceptual Loss. Prior research [15, 14, 49] has confirmed the effectiveness of perceptual loss in various domains. Notably, recent studies [26, 49] has found that the LPIPS loss [65] is crucial for obtaining few-step DMs with high sample quality. However, a notable drawback is that the LPIPS loss is computed in the data space, which is expensive. In contrast, the popular SD operate in the latent space to reduce computational demands. Hence, using LPIPS loss in training latent DMs is pretty expensive, as it requires not only the computation of LPIPS loss in the data space but also an additional decoding operation. Realizing that the pre-trained SD can serve as an effective feature extractor [58], we suggest using pre-trained SD to perform latent perceptual loss. However, SD is a UNet, whose final layer predicts epsilon with dimensions identical to the data. Hence we propose using the bottleneck layer of the UNet for computing:
@@
Latent Discriminator. Training GANs for the text-to-image on large-scale datasets faces serious challenges. Specifically, unlike unconditional generation, the discriminator for the text-to-image task should justify based on both image quality and text alignment. This challenge is more obvious during the initial stage of training. To address this issue, previous pure GANs [18] for text-to-image propose complex learning objectives and require expensive costs for training. The learning of GANs has been shown to benefit from using a pre-trained network as the discriminator. As discussed above, the pre-trained SD has learned representative features. Hence, we suggest applying the pre-trained SD for constructing the latent discriminator. Similar to latent perceptual loss, we only use half UNet for the discriminator followed by a simple predict head. The advantages of the proposed strategy are twofold: 1) we use the informative pre-trained network as initialization; 2) the discriminator is defined over latent space, which is computationally efficient. Unlike previous work [44] that defines a discriminator over data space which required decode latent and backward from decoder, yielding expensive computational cost. By applying the Latent Discriminator, we observe a stable training process with fast convergence. Note that, expensive R1 regularization is not needed in our approach.
Decoupled Scheduler. We find that the optimal scheduler for performing consistency loss and adversarial loss are not identical. Hence, to better unleash the capability of pre-trained DMs, we propose using decoupled schedulers to construct the following loss in fine-tuning pre-trained DMs:
where , , and we let and in experiments.
@@ 5.2 Fixing the Noise Scheduler
A common issue in DMs is that the final corrupted samples are not pure noise. For example, the noise scheduler used by the SD makes the corrupted samples at the final timestep as: , where the terminal Signal-to-noise ratio (SNR) is , effectively creates a gap between training and inference. Previous work [23] has only observed that this makes DMs unable to produce pure black or white images. However, we find that this issue yields serious problems in one-step generation. As shown in Fig. 2, there are notable artifacts in one-step generation if we directly sample noise from standard Gaussian distribution. The reason may be that in multi-step generation, the gap can be gradually fixed in sampling, while one-step generation reflects the gap more directly. To address this issue, we provide two simple yet effective solutions.
Informative Prior Initialization (IPI). The non-zero terminal SNR issue is similar to prior hole issues in VAEs [22, 3, 21]. Therefore, we can use informative prior instead of non-informative prior to effectively address this issue. For simplicity, we adopt a learnable Gaussian Distribution , whose optimal formulation is given below: , where and can be efficiently estimated by finite samples, and follows standard Gaussian distribution. As shown in Fig. 2, the artifacts in one-step generation are immediately removed after applying IPI. We note that the performance is achieved with minimal adjustment, enabling the possibility of developing one-step text-to-image generation by LoRA fine-tuning.
Quick Adaption to V-prediction and Zero Terminal SNR. The IPI suffers from numerical instability when terminal SNR is pretty low. As shown in Fig. 3, we fine-tune PixArt- [6] whose terminal SNR is 4e-5 with introduced technique and -prediction, failing in one-step generation. Following [23], we suggest switching to v-prediction [41] and zero terminal SNR. However, we found that direct transition convergences slowly (see Appendix F). This is unacceptable in solving large-scale text-to-image with limited computational resources. To address this, we propose a quick adapt-stage:
@@
We note that this adapt-stage converges rapidly, typically requiring only 1k iterations to initialize training YOSO. This enables a quick adaption to v-prediction and zero terminal SNR from pre-trained -prediction DMs, thereby comprehensively mitigating the non-zero terminal SNR issue in the noise scheduler.

### v5

@@ Abstract
Recently, some works have tried to combine diffusion and Generative Adversarial Networks (GANs) to alleviate the computational cost of the iterative denoising inference in Diffusion Models (DMs). However, existing works in this line suffer from either training instability and mode collapse or subpar one-step generation learning efficiency. To address these issues, we introduce YOSO, a novel generative model designed for rapid, scalable, and high-fidelity one-step image synthesis with high training stability and mode coverage. Specifically, we smooth the adversarial divergence by the denoising generator itself, performing self-cooperative learning. We show that our method can serve as a one-step generation model training from scratch with competitive performance. Moreover, we extend our YOSO to one-step text-to-image generation based on pre-trained models by several effective training techniques (i.e., latent perceptual loss and latent discriminator for efficient training along with the latent DMs; the informative prior initialization (IPI), and the quick adaption stage for fixing the flawed noise scheduler). Experimental results show that YOSO achieves the state-of-the-art one-step generation performance even with Low-Rank Adaptation (LoRA) fine-tuning. In particular, we show that the YOSO-PixArt- can generate images in one step trained on 512 resolution, with the capability of adapting to 1024 resolution without extra explicit training, requiring only ~10 A800 days for fine-tuning. Our code is provided at https://github.com/Luo-Yihong/YOSO.
@@ 1 Introduction
Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) have recently emerged as a powerful class of generative models, demonstrating state-of-the-art results in many generative modeling tasks, such as text-to-image (Rombach et al., 2022; Xu et al., 2023c; Chen et al., 2024; Feng et al., 2023), text-to-video (Blattmann et al., 2023; Hong et al., 2022), image editing (Hertz et al., 2022; Brooks et al., 2023; Meng et al., 2022) and controlled generation (Zhang et al., 2023; Mou et al., 2023). However, the generation process of DMs requires iterative denoising, leading to slow generation speed. Coupled with the intensive computational requirements of large-scale DMs, they constitute a substantial barrier to their practical application and wider adoption.
@@
Sampling from DMs can be regarded as solving a probability flow ordinary differential equation (PF-ODE) (Song et al., 2021). Some previous works (Song et al., 2020; Lu et al., 2022a; b; Bao et al., 2022) focus on developing advanced ODE-solvers, for reducing the sampling steps. However, they still require 20+ steps to achieve high-quality generation. Another line is distilling from pre-trained PF-ODEs (Song et al., 2023; Liu et al., 2023; Luo et al., 2023a; b; c; Salimans & Ho, 2022), aiming to predict multi-step solution of PF-ODE solver by one step. Existing methods (Luo et al., 2023a; b) can achieve reasonable sample quality with 4+ steps. However, it is still challenging to generate high-quality samples with one step.
@@
In contrast, generative adversarial networks (GANs) (Goodfellow et al., 2014; Radford et al., 2016) are naturally built on one-step generation with fast sampling speed. However, it is hard to extend GANs on large-scale datasets due to training challenges (Sauer et al., 2022; Kang et al., 2023), resulting in worse sample quality compared to DMs (Sauer et al., 2023b; Kang et al., 2023). In this work, we present a novel approach to combine diffusion process and GANs. The key to achieving access is finding a way to smooth the adversarial divergence for stabilize the training while maintaining effective one-step learning. Previous works (Xiao et al., 2022; Xu et al., 2023b; c; Sauer et al., 2023c) have developed some variants for combining diffusion and GANs. However, existing works either directly perform adversarial divergence against real data without smoothing strategy which suffers from unstable training and mode collapse, or rely on adding noise to smooth the adversarial divergence to stabilize the training which delivers less effective one-step generation learning. To enjoy the best of both worlds, we propose to smooth the adversarial divergence by denoising the generator itself. In particular, we regard the one-step denoising generation based on less corrupted samples as ground truth and regard the one-step denoising generation based on more corrupted samples as student distribution to perform adversarial divergence. The approach can not only naturally reduce the distance between target distribution and student distribution to stabilize the training, but also form effective one-step learning on clean samples. The learning process can be viewed as a self-cooperative process (Xie et al., 2018), as the generator learns from itself. Such an innovative design enables stable training and effective learning for one-step generation. Hence, we name our model YOSO, short for You Only Sample Once.
Moreover, we extend our YOSO to one-step text-to-image generation based on pre-trained models and introduce several effective training techniques (i.e., latent perceptual loss and latent discriminator for efficient training along with the latent DMs; the informative prior initialization (IPI), and the quick adaption stage for fixing the noise scheduler). Thanks to our effective designs, we can efficiently and effectively fine-tune existing pre-trained text-to-image DMs (i.e., stable diffusion (Rombach et al., 2022) and PixArt- (Chen et al., 2024)) for high-quality one-step generation (see Fig. 1). Furthermore, we are the pioneers in supporting Low-Rank Adaptation (LoRA) (Hu et al., 2022) fine-tuning in one-step text-to-image generation for enhanced efficiency, delivering state-of-the-art performance.
@@
We introduce YOSO, a novel generative model that can generate high-quality images with one-step inference, with a stable training process and good mode coverage.
We further scale up YOSO by several principled and effective proposed training techniques that enable low-resource fine-tuning of pre-trained text-to-image DMs for one-step text-to-image, requiring only ~10 A800 days.
We conduct extensive experiments to demonstrate the effectiveness of the proposed YOSO, including image generation from scratch, text-to-image generation fine-tuning, compatibility with existing image customization and image controllable modules.
@@ 2 Background
Diffusion models. Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020) define a forward process that gradually transforms samples from data distribution to Gaussian distribution by adding noise in steps with variance schedule : . The corrupted samples can be directly obtained by , where and . The parameterized reversed diffusion process is defined to gradually denoise: . The model can be trained by minimizing the negative ELBO (Ho et al., 2020; Kingma et al., 2021): where is Gaussian posterior derived in (Ho et al., 2020). A key assumption in diffusion is that the denoising step size from to is sufficiently small. This assumption ensures the true follows Gaussian distribution, enabling the effectiveness of modeling with Gaussian distribution.
Diffusion-GAN hybrids. An issue in DMs is that the true does not follow Gaussian distribution when the denoising step size is not sufficiently small. Therefore, in order to enable large denoising step size, Diffusion GANs (Xiao et al., 2022) propose to minimize the adversarial divergence between model and : , where and is imposed by a GAN generator. The capability of a GAN-based formulation enables much larger denoising step sizes (i.e., 4 steps).
@@ 3 Method: Self-Cooperative Diffusion GANs
A key issue in Diffusion-GAN hybrid models (Xiao et al., 2022; Xu et al., 2023b; c) is that they match the generator distribution with the corrupted data distribution. This formulation only indirectly learns the and , which are the distributions used for one-step generation, making the learning process less effective.
@@ 3.1 Our Design
where and . The optimization objective is constructed by combining adversarial divergence and KL divergence. Specifically, the adversarial divergence focuses on matching over the distribution level, ensuring the generation quality, and the KL divergence focuses on matching over point level, ensuring the mode coverage.
However, it is hard to directly learn the adversarial divergence over clean data distribution, akin to the challenges with GAN training. To tackle these challenges, previous Diffusion GANs (Xiao et al., 2022; Xu et al., 2023b) have pivoted towards learning the adversarial divergence over the corrupted data distribution. Unfortunately, as analyzed before, such an approach fails to directly match , curtailing the efficacy of one-step generation. Moreover, it also compels the discriminator to fit different levels of noise, leading to limited capability.
@@
Hence given the generator fixed, if we increase the information in , it is supposed that we can get a better distribution. In other words, it is highly likely that is superior to , where . Motivated by the cooperative approach (Xie et al., 2018; 2021; 2022; Hill et al., 2022) which uses an MCMC revised version of model distribution to learn the generator, we suggest to use as ground truth to learn , which constructs the following training objective:
where denotes stop-gradient operator and the second term is named as cooperative adversarial loss. This training objective can be regarded as a self-cooperative approach, since the ‘revised’ samples from and samples from are generated by the same network . Note that we only replace data distribution as in the adversarial divergence for smoothing the learning objective, as recent work (Luo et al., 2023d) has found that it is beneficial to learn the generator with a mix of real data and revised data.
We briefly verify the theoretical rationality of the cooperative adversarial divergence below.
Proposition 1
The optimal solution of the cooperative adversarial loss reaches .
This proposition tells that the proposed cooperative adversarial loss can recover the real data distribution when the network’s capability is sufficiently large, demonstrating the theoretical rationality of the proposed objective. See proof in the Appendix C.
@@
where is the discriminator network. We find that the self-cooperative approach is connected with Consistency Training (Song et al., 2023). However, Consistency Training considers the as an approximated ODE solution of to perform a point-to-point match. In contrast, our proposed objective matches and in marginal distribution level, which avoids the approximated error of ODE.
To further ensure the mode cover of the proposed model, we can add consistency loss to our objective as regularization, which constructs the following loss:
where and are pre-defined hyper-parameters.
4 Try It On CIFAR-10 Before Scaling Up For Saving Money!
In this section, we evaluate the performance of the proposed YOSO on CIFAR-10 (Yu et al., 2015) to verify its effectiveness under both training from scratch and fine-tuning settings.
4.1 training strategies
Before starting training, we introduce some effective training strategies in following for taming YOSO.
Decoupled scheduler. We find that the optimal scheduler for performing consistency loss and adversarial loss are not identical. This is due to the cooperative adversarial loss not involving the approximated error regarding the timestep skips, enabling a substantial skip to maximize its efficacy. In contrast, the consistency loss is susceptible to approximated error from timestep skips, necessitating a more conservative skip to preserve its effectiveness. Hence, to better unleash the capability of each loss, we propose using decoupled schedulers to construct our final training objective:
where , , and we let and in experiments.
Model | FID | NFE |
Distillation or Fine-tuning | ||
YOSO | 1.81 | 1 |
Tract (Berthelot et al., 2023) | 3.78 | 1 |
Diff-Instruct (Luo et al., 2023c) | 4.53 | 1 |
DMD (Yin et al., 2023) | 2.66 | 1 |
CTM (Kim et al., 2024) | 1.98 | 1 |
SiD (Zhou et al., 2024) | 1.92 | 1 |
Training from scratch | ||
YOSO | 2.26 | 1 |
DDGANs (Xiao et al., 2022) | 3.75 | 4 |
CT (Song et al., 2023) | 8.70 | 1 |
iCT (Song & Dhariwal, 2024) | 2.83 | 1 |
DDPM (Ho et al., 2020) | 3.21 | 1000 |
EDM (Karras et al., 2022) | 2.04 | 36 |
DDIM (Song et al., 2020) | 4.67 | 50 |
StyleGAN2 (Karras et al., 2019) | 2.92 | 1 |
Annealing strategy. Since our target is to obtain a powerful one-step denoised generation model, however, as the KL loss and consistency loss will enforce point-level matching, this may hurt performance in cases where the model capacity is insufficient. Therefore, we suggest reducing the weight of these two losses to zero as training progresses. Specifically, we let for decreasing the weight to zero with times, where is the current iterations and is the total training iterations.
4.2 Empirical Evaluation
Experiment Setting. We use the EDM (Karras et al., 2022) architecture for both UNet and discriminator. For the evaluation metric, we choose the Fréchet inception distance (FID) (Heusel et al., 2017). We train the proposed model on the CIFAR-10 dataset (Yu et al., 2015). We only employ MSE to compute image distance.
Results. The quantitative results are shown in Tab. 1. We observe that our method provides state-of-the-art performance with only a one-step manner on both training from scratch and fine-tuning settings, outperforming previous existing accelerated DMs, and GANs. In particular, our YOSO in fine-tuning EDM (Karras et al., 2022) even delivers better performance compared to EDM itself with ODE sampling. This is because our YOSO performs fine-tuning instead of distillation, the optimal solution for our YOSO is data distribution instead of teacher distribution.
@@ 5.1 Using pre-trained models for Training
Latent Perceptual Loss. Prior research (Hou et al., 2017; Hoshen et al., 2019; Song et al., 2023) has confirmed the effectiveness of perceptual loss in various domains. Notably, recent studies (Liu et al., 2023; Song et al., 2023) have found that the LPIPS loss (Zhang et al., 2018) is crucial for obtaining few-step DMs with high sample quality. However, a notable drawback is that the LPIPS loss is computed in the data space, which is expensive. In contrast, the popular SD operates in the latent space to reduce computational demands. Hence, using LPIPS loss in training latent DMs is pretty expensive, as it requires not only the computation of LPIPS loss in the data space but also an additional decoding operation. Realizing that the pre-trained SD can serve as an effective feature extractor (Xu et al., 2023a), we suggest using pre-trained SD to perform latent perceptual loss. However, SD is a UNet, whose final layer predicts epsilon with dimensions identical to the data. Hence we propose using the bottleneck layer of the UNet for computing:
@@
Latent Discriminator. Training GANs for the text-to-image on large-scale datasets faces serious challenges. Specifically, unlike unconditional generation, the discriminator for the text-to-image task should justify based on both image quality and text alignment. This challenge is more obvious during the initial stage of training. To address this issue, previous pure GANs (Kang et al., 2023) for text-to-image propose complex learning objectives and require expensive costs for training. The learning of GANs has been shown to benefit from using a pre-trained network as the discriminator. As discussed above, the pre-trained SD has learned representative features. Hence, we suggest applying the pre-trained SD for constructing the latent discriminator. Similar to latent perceptual loss, we only use half UNet for the discriminator followed by a simple predict head. The advantages of the proposed strategy are twofold: 1) we use the informative pre-trained network as initialization; 2) the discriminator is defined over latent space, which is computationally efficient. Unlike previous work (Sauer et al., 2023c) that defines a discriminator over data space which required decode latent and backward from decoder, yielding expensive computational cost. By applying the Latent Discriminator, we observe a stable training process with fast convergence.
@@ 5.2 Fixing the Noise Scheduler
A common issue in DMs is that the final corrupted samples are not pure noise. For example, the noise scheduler used by the SD makes the corrupted samples at the final timestep as: , where the terminal Signal-to-noise ratio (SNR) is , effectively creates a gap between training and inference. Previous work (Lin et al., 2024a) has only observed that this makes DMs unable to produce pure black or white images. However, we find that this issue yields serious problems in one-step generation. As shown in Fig. 2, there are notable artifacts in one-step generation if we directly sample noise from standard Gaussian. The reason may be that in multi-step generation, the gap can be gradually fixed in sampling, while one-step generation reflects the gap more directly. To address this issue, we provide two simple yet effective solutions.
Informative Prior Initialization (IPI). The non-zero terminal SNR issue is similar to prior hole issues in VAEs (Klushyn et al., 2019; Bauer & Mnih, 2019; Kingma et al., 2016). Therefore, we can use informative prior instead of non-informative prior to effectively address this issue. For simplicity, we adopt a learnable Gaussian Distribution , whose optimal formulation is given below: , where and can be efficiently estimated by finite samples, and follows standard Gaussian distribution. As shown in Fig. 2, the artifacts in one-step generation are immediately removed after applying IPI. We note that the performance is achieved with minimal adjustment, enabling the possibility of developing one-step text-to-image generation by LoRA fine-tuning.
Quick Adaption to V-prediction and Zero Terminal SNR. The IPI suffers from numerical instability when terminal SNR is pretty low. As shown in Fig. 3, we fine-tune PixArt- (Chen et al., 2024) whose terminal SNR is 4e-5 with introduced technique and -prediction, failing in one-step generation. Following (Lin et al., 2024a), we suggest switching to v-prediction (Salimans & Ho, 2022) and zero terminal SNR. However, we found that direct transition convergences slowly (see Appendix F). This is unacceptable in solving large-scale text-to-image with limited computational resources. To address this, we propose a quick adapt-stage:
@@
We note that this adapt stage converges rapidly, typically requiring only 1k iterations to initialize training YOSO. This enables a quick adaption to v-prediction and zero terminal SNR from pre-trained -prediction DMs, thereby comprehensively mitigating the non-zero terminal SNR issue in the noise scheduler.